The following information is related to Arabian Food manufacturing company as of 31st

December 2020

 The company has 50 million common stock shares worth of 400 million dirhams.

 The company’s stocks are currently trading at 20 dirhams per share.

 The company recently paid 3 dirhams per share and expected to grow at 5% annually.

 The company has issued preference shares for a value of 50 million dirhams consisting

of 5 million preference shares and the annual dividend per share is 10 dirhams per share.

 The last traded price of a preference share was 125 dirhams per share.

 The average cost of debt for the company’s long-term debts worth of 150 million

dirhams is 12% per annum.

 The company’s marginal tax rate is 30%.


Required?

(a) Calculate the costs of each sources of capital.

(b) What is the weighted average cost of capital of the company?


1
Expert's answer
2021-04-19T18:48:15-0400

(a) The costs of each sources of capital are:

Cost of equity is:

Re = DPS/MPS + r = 3/20 + 0.05 = 0.2.

Cost of debt is 0.12.

Cost of preference shares is: Rp = D/P = 10/125 = 0.08.

(b) What is the weighted average cost of capital of the company?

"WACC = (E\/(E + D)\u00d7Re) + ((D\/(E + D)\u00d7Rd) \u00d7(1 \u2013 T)) + P\/(E + D)\u00d7Rp = 400,000\/(400,000+150,000 + 50,000)\u00d70.2+150,000\/600,000\u00d70.12\u00d7(1-0.3) + 50,000\/600,000\u00d70.08=0.161."